OpenStudy (anonymous):

Expand (1 + y)^8 in ascending powers of y, up to y^3. In the expansion (1 + x + kx)^8 in ascending powers of x, the coefficient of x^3 is zero. Find the constant k. Much thanks in advance! :)

OpenStudy (anonymous):

\((1+y)^8=1+8y+28y^2+56y^3+...\) coefficients come from \(\binom{8}{k}\) for \(k=0,1,2,3\)

OpenStudy (anonymous):

so for example the coefficient of the \(y^2\) term is \(\binom{8}{2}=\frac{8\times 7}{2}=4\times 7=28\)

OpenStudy (anonymous):

was this really posted ten hours ago?

OpenStudy (anonymous):

@satellite73 I'm sorry for replying to a week-old post. D: I haven't had the time to go back to it lately, I'm really really sorry. Thank you so much for your help!